Two families are saving up for a summer trip. The Smith family originally has $1000 and will save $50 every week. The Jones family originally has $500 and will save $75 every week.
After how many weeks will the two families have saved up the same amount of money?

Answer by addingup(3677)

1000 + 50x = 500 + 75x
500 = 25x
x = 500/25 = 20
In 20 weeks the two families will have saved up the same anount of money